# The Cone Restriction: An old Approach Revisited

### v4

@@ 2.5 Transverse Equidistribution Estimates
If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then the Lemma 2.4 gives
@@
[…]
@@
Suppose $A$ is large enough. By the definition of $k$-board norm, and lemma 2.24,
@@ 2.6 Polynomial Partitioning
Suppose we have a polynomial $P$ described in the theorem 2.27. Let ${\delta}^{\prime}>0$ be a small number. Let $\mathcal{O}$ denote the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)$. By pigeonholing, we see that at least one of two cases hold:
@@
(1).$\mathcal{O}\sim d^{m}$, and each $O\in\mathcal{O}$ is contained in a ball of radius $r/2$.
@@
Now we apply the theorem 2.28 to the $k$-board norm estimates:
@@
#$\#\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
@@ 3 Finding Polynomial Structure
In this section, we reformulate the induction arguments in [4] as a recursive process by combining the ideas from [12, 13]. The argument contains two different algorithms. In [alg 1] we apply the iterated polynomial partition until the tangent term dominates. This algorithm passes from an $m$-dimensional to an $\left(m-1\right)$-dimensional situation. This algorithm corresponds to the induction on the scale. [alg 2] is established by repeated application of [alg 1]. The iteration ends when we arrive in the minimal dimensional case. The lemma 2.16 is used to establish the bound. This algorithm corresponds to the induction on the dimension.
@@ 3.1 The first algorithm
$\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning. Let #$\#\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and #$\#\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
@@
$\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$. Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)$.
@@
$\bullet$ A large integer $d$ which depends only on the admissible parameters and ${\mathrm{deg}\boldsymbol{\mathrm{Z}}\ }$.
@@ 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
@@ 1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
2. For all $O\in\mathcal{O}$, $O$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
@@
By lemma 2.29, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus the lemma 2.1 gives
@@
By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with #$\#{\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
@@
Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\#\mathcal{B}\ \sim\ 1$ such that
@@
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in the property 4. Combine (3.5) and (3.7),
@@
Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Then, the definition of the parameters gives
#$\#\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
@@
Since #$\#{\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, the lemma 2.1 gives
@@
Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, the lemma 2.1 and corollary 2.7 gives
@@
#$\#\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
@@ 3.2 The second algorithm
$\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain. Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and bounded degree. Also, for $l\leq s\leq t\leq n$:
@@
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$. Let #$\#\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and #$\#\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote #$\#\mathrm{a}\left(J\right)$ and #$\#\mathrm{c}\left(J\right)$ output by [alg 1] respectively. Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively. Let
[…]
@@
By lemma 2.20, one has $m\geq k$. If $m<k$, then lemma 2.20 implies
@@
Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20). Property 3 follows from (LABEL:r3.26_) and the repeated application of (3.21). Property 1 follows from (3.19), (3.37), (3.24), (3.35), and lemma 2.18. See [12] for details. Here we verify property 4. Apply (3.23) repeatedly, one has
@@ 3.3 Establish the $k$-broad estimates
This gives the main result of Ou-Wang [4]. See [12] for details.
@@
In this section we adapt the method provided in [13] to prove theorem 2.22. Let $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ and $A$ is a large integer. Without loss of generality, we assume that $f$ satisfies (3.27). By applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described as Section 3.2.
First, we improve (3.42) by using lemma 2.16, the nested polynomial Wolff axioms.
@@
Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]$, lemma 2.16 implies
@@
Then lemma 4.1 follows from (4.1) and (4.2). ∎
By combining (3.32), (3.33), and lemma 4.1, for $m\leq l<n-1$,
@@
$=\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
@@
By taking the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$, we have

### v5

@@ 2.5 Transverse Equidistribution Estimates
If ${\left\|Eg\right\|}^{2}_{L^{2}\left(B\left(y,r\right)\right)}$ doesn’t vanish, then Lemma 2.4 gives
@@
$V^{+}:=\left\{\sum^{m}_{i=1}{a_{i}v^{+}_{i}}:a_{i}\in\mathbb{R}\mathrm{,}v^{+}_{i}=\left(x_{i,1},x_{i,2},\dots,x_{i,n}\right),\ \mathrm{f}\mathrm{or\ }1\leq i\leq m\right\}$
@@
Suppose $A$ is large enough. By the definition of $k$-board norm and Lemma 2.25,
@@ 2.6 Polynomial Partitioning
Suppose we have a polynomial $P$ described in Theorem 2.28. Let ${\delta}^{\prime}>0$ be a small number. Let $\mathcal{O}$ denote the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}^{\prime}}}Z\left(P\right)$. By pigeonholing, we see that at least one of two cases hold:
@@
[…]
@@
Now we apply Theorem 2.29 to the $k$-board norm estimates:
@@
$\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
@@ 3 Finding Polynomial Structure
In this section, we reformulate the induction arguments in [15] as a recursive process by combining the ideas from [11, 12]. The argument contains two different algorithms. In [alg 1] we apply the iterated polynomial partition until the tangent term dominates. This algorithm passes from an $m$-dimensional to an $\left(m-1\right)$-dimensional situation. This algorithm corresponds to the induction on the scale. [alg 2] is established by repeated application of [alg 1]. The iteration ends when we arrive in the minimal dimensional case. Corollary 2.17, the nested polynomial Wolff axioms, is used to establish the bound. This algorithm corresponds to the induction on the dimension.
@@ 3.1 The first algorithm
$\bullet$ History: A word ${\hslash}_{j}$ of length $j$ in the alphabet $\left\{\mathrm{a,c}\right\}$, which records how the cell $O_{j}\in{\mathcal{O}}_{j}$ were constructed by the iterated polynomial partitioning. Let $\mathrm{a}\left(j\right)$ denote the number of ‘a’ in ${\hslash}_{j}$, and $\mathrm{c}\left(j\right)$ denote the number of ‘c’ in ${\hslash}_{j}$.
@@
$\bullet$ Cells: A family of subsets ${\mathcal{O}}_{j}$ of ${\mathbb{R}}^{n}$. Each cell $O_{j}\in{\mathcal{O}}_{j}$ is contained in some ${\rho}_{j}$-ball $B_{O_{j}}=B\left(y_{O_{j}},{\rho}_{j}\right)\$.
@@
$\bullet$ A large integer $d$ which depends only on the admissible parameters and $\mathrm{deg}\boldsymbol{\mathrm{Z}}\$}.
@@ 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j}\right]$ and $O_{j}\in{\mathcal{O}}_{j}$,
Define ${\hslash}_{j+1}$ by appending the $\mathrm{c}$ to ${\hslash}_{j}$. Now
@@ 1. $\mathcal{O}\left(O_{j}\right)\ \sim\ d^{m}$
2. For all $\in\mathcal{O}, OO$ is included in a ${\rho}_{j+1}$-ball $B_{O}:=B\left(y_{O},{\rho}_{j+1}\right)$.
@@
By Lemma 2.30, each $T^{l}_{\theta,v}$ intersects $O\left(d\right)$ cells, and thus Lemma 2.1 gives
@@
By pigeonholing, there exists a refinement ${\mathcal{O}}^{\prime}\left(O_{j}\right)\subseteq\mathcal{O}\left(O_{j}\right)$ with ${\mathcal{O}}^{\prime}\ \sim\ d^{m}$ such that
@@
Since ${\rho}_{j+1}\sim{\rho}_{j}$, there exists $\mathcal{B}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ with # $\mathcal{B}\ \sim\ 1$ such that
@@
for some $\left(\zeta,w,L\right)\in\mathcal{T}$. Here $C$ is the absolute constant that appears in Property 4. Combine (3.5) and (3.7),
@@
Define ${\hslash}_{j+1}$ by appending the $\mathrm{a}$ to ${\hslash}_{j}$. Now
$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
@@
Since ${\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, Lemma 2.1 gives
@@
Also, for $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$ and $S=\left(Y,B\right)\in\mathcal{S}$ with $B\in{\mathcal{B}}_{O_{j}}$, Lemma 2.1 and Corollary 2.7 gives
@@
$\mathrm{c}\left(J\right)\lesssim{\mathrm{log}r\ },\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}{\mathrm{log}{\delta}^{-1}_{m-1}\ }$ (3.18)
@@ 3.2 The second algorithm
$\bullet$ For $m\leq l\leq n$ a family ${\overrightarrow{\mathcal{S}}}_{l}$ of level $\left(n-l+1\right)$ multigrain. Each ${\overrightarrow{S}}_{l}=\left(\left(Z_{l},B_{r_{l}}\right),\dots,\left(Z_{n},B_{r_{n}}\right)\right)\in{\overrightarrow{\mathcal{S}}}_{l}$ has a multiscale ${\overrightarrow{r}}_{l}=\left(r_{l},\dots,r_{n}\right)$ and a bounded degree. Also, for $l\leq s\leq t\leq n$:
@@
For each ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}$ we apply [alg 1] to the $k$-board norm ${\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}$. Let $\mathrm{a}\left[{\overrightarrow{S}}_{l}\right]$ and $\mathrm{c}\left[{\overrightarrow{S}}_{l}\right]$ denote $\mathrm{a}\left(J\right)$ and $\mathrm{c}\left(J\right)$ output by [alg 1] respectively. Let $A\left[{\overrightarrow{S}}_{l}\right]$ and $r\left[{\overrightarrow{S}}_{l}\right]$ denote $A_{J}$ and $r_{J}$ output by [alg 1] respectively. Let
${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{iny}\mathrm{]}\right\}$
${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{ang}\mathrm{]}\right\}$
@@
If $m<k$, then Lemma 2.21 implies
@@
Verify the properties Property 2 follows from (3.25) and the repeated application of (3.20). Property 3 follows from (3.26) and the repeated application of (3.21). Property 1 follows from (3.19), (3.37), (3.24), (3.35), and Lemma 2.19. See [11] for details. Here we verify Property 4. Apply (3.23) repeatedly, one has
@@ 3.3 Establish the $k$-broad estimates
This gives Theorem 1.3. See [11] for details.
@@
In this section we adapt the method provided in [12] to prove Theorem 2.23. Let $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ and $A$ is a large integer. Without loss of generality, we assume that $f$ satisfies (3.27). Applying [alg 2] to decompose the $k$-broad norm ${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}$, we get the outputs described in Section 3.2.
First, we improve (3.42) by Corollary 2.17, the nested polynomial Wolff axioms.
@@
Since $f^{\#}_{{\overrightarrow{S}}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $\left(\theta,v,l\right)$ with $\theta\in\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right$], Corollary 2.17 implies
@@
Now Lemma (4.1) follows from (4.1) and (4.2). ∎
Combining (3.32), (3.33), and Lemma 4.1, we see that for $m\leq l<n-1$,
@@
$\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right)\left(\prod^{n-1}_{j=m-1}{D^{\delta}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.3)
@@
Take the geometric mean of (3.41) and (4.3) with weights $\gamma_{i}$,